# Reinforcement Learning Teachers of Test Time Scaling

### v2

@@ Appendix D Extended analysis
Postprocessed R1 reasoning distillation example Masked distillation input <|im_start|>system <|im_start|>system Your role as an assistant involves thoroughly exploring questions... <|im_start|>user Generate an executable Python function generated from the given prompt... <|im_end|> <|im_start|>assistant Distillation reasoning trace <|begin_of_thought|> Okay, let’s see. The problem is to find the minimum number of sheets needed to print N pages double-sided. Each sheet can hold two pages. So, I need to figure out how to calculate this based on N. Hmm, let’s think about some examples. Like the sample input is 5, output is 3. Because 5 divided by 2 is 2.5, which rounds up to 3. So maybe the general formula is to divide N by 2 and round up. Wait, but wait. Let me check another example. Suppose N is 4. Then 4/2 is exactly 2 sheets. That makes sense. If N is 1, then 1/2 is 0.5, which rounds up to 1. So that’s correct. So the formula is simply ceiling(N / 2). But how to compute that in Python. In Python, if you do integer division with //, but when you need to round up, you can do (N + 1) // 2. Let’s see. For N=5: (5 +1)//2 =6//2=3. Correct. For N=4: (4+1)/2=5/2=2.5, integer division gives 2. Wait, no. Wait, 5//2 in Python is 2. So (4+1)//2 is 5//2=2. Which is correct. For N=1: (1+1)//2=1. Correct. So yes, the formula is (N +1) // 2. So the function is straightforward. Take the input N, compute (N +1)//2, and output it. Testing the sample input: input 5. (5+1)//2=6//2=3. Correct. Another test case: N=1 → 1. N=2 → 1. N=3 →2. All correct. So the code would be something like: Read N from input, compute (N+1)//2, print the result. So the Python code would be: n = int(input()) print( (n +1) // 2 ) But wait, perhaps there’s an edge case. Let’s see. For N=0? But according to constraints, N is at least 1, so no problem. Yes, so that’s the solution. The code is simple. Just read N, calculate (N +1)//2, print. <|end_of_thought|> Distillation solution <|begin_of_solution|> To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. 3. **Output**: The result is printed, which is the minimum number of sheets needed. This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>
@@
Raw 7B RLT distillation example Masked distillation input <|im_start|>system <|im_start|>system Your role as an assistant involves thoroughly exploring questions... <|im_start|>user Generate an executable Python function generated from the given prompt... <|im_end|> <|im_start|>assistant Distillation reasoning trace <|begin_of_thought|> Okay, let’s see. The problem is to find the minimum number of sheets needed to print N pages double-sided. Each sheet can hold two pages. So, I need to figure out how to calculate this based on N. Hmm, let’s think about some examples. Like the sample input is 5, output is 3. Because 5 divided by 2 is 2.5, which rounds up to 3. So maybe the general formula is to divide N by 2 and round up. Wait, but wait. Let me check another example. Suppose N is 4. Then 4/2 is exactly 2 sheets. That makes sense. If N is 1, then 1/2 is 0.5, which rounds up to 1. So that’s correct. So the formula is simply ceiling(N / 2). But how to compute that in Python. In Python, if you do integer division with //, but when you need to round up, you can do (N + 1) // 2. Let’s see. For N=5: (5 +1)//2 =6//2=3. Correct. For N=4: (4+1)/2=5/2=2.5, integer division gives 2. Wait, no. Wait, 5//2 in Python is 2. So (4+1)//2 is 5//2=2. Which is correct. For N=1: (1+1)//2=1. Correct. So yes, the formula is (N +1) // 2. So the function is straightforward. Take the input N, compute (N +1)//2, and output it. Testing the sample input: input 5. (5+1)//2=6//2=3. Correct. Another test case: N=1 → 1. N=2 → 1. N=3 →2. All correct. So the code would be something like: Read N from input, compute (N+1)//2, print the result. So the Python code would be: n = int(input()) print( (n +1) // 2 ) But wait, perhaps there’s an edge case. Let’s see. For N=0? But according to constraints, N is at least 1, so no problem. Yes, so that’s the solution. The code is simple. Just read N, calculate (N +1)//2, print. <|end_of_thought|> Distillation solution <|begin_of_solution|> To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. 3. **Output**: The result is printed, which is the minimum number of sheets needed. This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>
@@
Figure 10: Full formatted distillation data sample using the raw reasoning trace from our 7B RLT model.
@@
Thought tokens KL reward ablation distillation example Masked distillation input <|im_start|>system <|im_start|>system Your role as an assistant involves thoroughly exploring questions... <|im_start|>user Generate an executable Python function generated from the given prompt... <|im_end|> <|im_start|>assistant Distillation reasoning trace <|begin_of_thought|> To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. 3. **Output**: The result is printed, which is the minimum number of sheets needed. This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_thought|> Distillation solution <|begin_of_solution|> To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. 3. **Output**: The result is printed, which is the minimum number of sheets needed. This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>
@@
Min/max reduction reward ablation distillation example Distillation reasoning trace <|begin_of_thought|> Okay, I need to solve this problem where Takahashi wants to print N pages double-sided, with each sheet holding two pages. The task is to find the minimum number of sheets required. Let’s think through this step by step. So, the problem is about packing pages into sheets. Each sheet can have two pages. But since it’s double-sided, maybe the way the pages are arranged affects how they fit. Wait, no, double-sided printing just means that two pages are printed on the same sheet, front and back. So for example, if I have page 1 on the front and page 2 on the back, that’s one sheet. But if there’s an odd number of pages, the last sheet will have only one page. For example, if N is 5, then the sample input shows 3 sheets. Let’s see: 1-2 on first sheet, 3-4 on second, 5 on third. So each sheet except possibly the last can have two pages. Hmm, so the approach is to calculate how many pairs we can make. For N pages, the number of pairs is N//2. But if N is odd, there’s an extra page. So the total number of sheets would be (N // 2) + (1 if N % Wait, but wait. Let’s check another example. If N is 4. Then 4//2 is 2, and 4% Another example: N=3. 3//2 is 1, 3% What about N=1? Then 1//2 is 0, 1% Wait, the problem says each sheet can hold two pages. So if there’s an odd number of pages, the last sheet has one page. But the problem’s sample input is 5, which works out to 3 sheets. So maybe when N is odd, the formula is (N+1)//2. Wait, 5+1 is 6, 6//2 is 3. Which is correct. For N=3, 3+1=4, 4//2=2. Correct. For N=1, 1+1=2, 2//2=1. So that’s correct. Wait, 1+1=2, 2//2=1. So the formula is (N + 1) // 2. Let’s test that. Wait, for N=1: (1+1)//2 = 1. Which is correct. For N=2: (2+1)/2=1.5, integer division is 1. Which is correct. For N=3: (3+1)/2=2. Correct. N=4: 5//2=2.5→2. Correct. N=5: 3. Correct. So this formula seems to work. Wait, but wait. Let’s see. For N=1, if we use (N+1)//2, that’s (1+1)//2=1. So yes. So the formula is (N +1)//2. But wait, in the problem statement, the sample input 5 gives output 3, which is (5+1)//2=3. So that’s correct. So the formula is (N +1)//2. Because adding 1 gives even number of pages, which can be divided by 2, and then integer division by 2 gives the correct number of sheets. Wait, but let me think again. For N=1, the formula gives 1. So that’s correct. For N=2, (2+1)//2=1.5→1. Which is correct. So yes, this formula works. Wait, but why would (N+1)//2 be the correct formula? Because when you have N pages, each sheet holds two pages, except possibly the last one. So the total number of sheets is the number of pairs (each pair is two pages) plus one if there’s an odd number. So the number of pairs is N//2, and if N is odd, there’s an extra sheet. So the total is (N//2) + (1 if N % Yes. Because for N even, (N+1)/2 is (N/2) +0.5 → but integer division would be N/2. For example, N=4: (4+1)/2=2.5→2. Which is correct. For N=5: (5+1)/2=3. So that’s correct. So the formula is (N +1) // 2. Wait, but wait. Let’s test N=1. (1+1)/2=1. Correct. N=2: (2+1)/2=1.5→1. Correct. N=3: (3+1)/2=2. Correct. N=4: 2. Correct. N=5:3. Correct. So this formula works. So the problem reduces to calculating (N + 1) // 2. But wait, the input is N, which is between 1 and 100. So the function needs to read N from standard input, compute (N+1)//2, and output that. Wait, but wait. Let’s take N=1. The sample input is 5, which gives output 3. For N=1, the output would be 1. But according to the problem statement, the minimal number of sheets. So for N=1, you need 1 sheet. Because even though you can’t fill a full sheet, you just have to use one sheet. So the formula works. So the code is straightforward. Read N, compute (N+1)//2, print that. Wait, but let me check another example. Let’s say N=6. (6+1)//2=3.5→3. Wait, 6//2=3. Correct. So that’s correct. Wait, but wait. Wait, no. Wait, (6+1)//2 is 7//2=3. Wait, 7 divided by 2 is 3.5, integer division is 3. So (6+1)//2 is 3.5 →3. So yes. So 6 pages would require 3 sheets. Each sheet has two pages. So 1-2, 3-4, 5-6. So three sheets. Correct. So the code is as simple as reading N, then compute the ceiling of N divided by 2. Which is (N +1)//2. So the Python function would be: Read N from input. Then compute (N +1) // 2. Print that result. Wait, but wait. Let’s test N=1 again. (1+1)//2=1. Correct. Yes. So the code is straightforward. So the code would be: Read N as integer from input. Then print (N + 1) // 2. But wait, wait. Let’s check N=1. (1+1)//2=1. Correct. For N=2: (2+1)//2=1. So 1 sheet. Which is correct. For N=3: 2 sheets. Correct. For N=4: 2. Correct. For N=5:3. Correct. For N=6:3. Correct. So the code is correct. Therefore, the Python function is: Read N from input, then compute (N +1) // 2 and print it. So the code would be: n = int(input()) print((n + 1) // 2) Wait, but wait. Let me think again. For N=1, the answer is 1. For N=2, 1. So that’s correct. So the code is correct. But wait, the problem says that N is the number of pages. So that’s correct. So the code is as simple as that. I think that’s it. So the answer is to compute (N +1) // 2. <|end_of_thought|> Distillation solution <|begin_of_solution|> To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. 3. **Output**: The result is printed, which is the minimum number of sheets needed. This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>

### v3

@@ Appendix D Extended analysis
Postprocessed R1 reasoning distillation example Masked distillation input <|im_start|>system <|im_start|>system Your role as an assistant involves thoroughly exploring questions... <|im_start|>user Generate an executable Python function generated from the given prompt... <|im_end|> <|im_start|>assistant Distillation reasoning trace <|begin_of_thought|> {internallinenumbers*} Okay, let’s see. The problem is to find the minimum number of sheets needed to print N pages double-sided. Each sheet can hold two pages. So, I need to figure out how to calculate this based on N. {internallinenumbers*} Hmm, let’s think about some examples. Like the sample input is 5, output is 3. Because 5 divided by 2 is 2.5, which rounds up to 3. So maybe the general formula is to divide N by 2 and round up. {internallinenumbers*} Wait, but wait. Let me check another example. Suppose N is 4. Then 4/2 is exactly 2 sheets. That makes sense. If N is 1, then 1/2 is 0.5, which rounds up to 1. So that’s correct. So the formula is simply ceiling(N / 2). But how to compute that in Python. {internallinenumbers*} In Python, if you do integer division with //, but when you need to round up, you can do (N + 1) // 2. Let’s see. For N=5: (5 +1)//2 =6//2=3. Correct. For N=4: (4+1)/2=5/2=2.5, integer division gives 2. Wait, no. Wait, 5//2 in Python is 2. So (4+1)//2 is 5//2=2. Which is correct. For N=1: (1+1)//2=1. Correct. So yes, the formula is (N +1) // 2. {internallinenumbers*} So the function is straightforward. Take the input N, compute (N +1)//2, and output it. Testing the sample input: input 5. (5+1)//2=6//2=3. Correct. Another test case: N=1 → 1. N=2 → 1. N=3 →2. All correct. So the code would be something like: Read N from input, compute (N+1)//2, print the result. So the Python code would be: n = int(input()) print( (n +1) // 2 ) {internallinenumbers*} But wait, perhaps there’s an edge case. Let’s see. For N=0? But according to constraints, N is at least 1, so no problem. {internallinenumbers*} Yes, so that’s the solution. The code is simple. Just read N, calculate (N +1)//2, print. <|end_of_thought|> Distillation solution <|begin_of_solution|> {internallinenumbers*} To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach {internallinenumbers*} The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation {internallinenumbers*} 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. {internallinenumbers*} 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. {internallinenumbers*} 3. **Output**: The result is printed, which is the minimum number of sheets needed. {internallinenumbers*} This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>
@@
Raw 7B RLT distillation example Masked distillation input <|im_start|>system <|im_start|>system Your role as an assistant involves thoroughly exploring questions... <|im_start|>user Generate an executable Python function generated from the given prompt... <|im_end|> <|im_start|>assistant Distillation reasoning trace <|begin_of_thought|> {internallinenumbers*} Okay, let’s see. The problem is to find the minimum number of sheets needed to print N pages double-sided. Each sheet can hold two pages. So, I need to figure out how to calculate this based on N. {internallinenumbers*} Hmm, let’s think about some examples. Like the sample input is 5, output is 3. Because 5 divided by 2 is 2.5, which rounds up to 3. So maybe the general formula is to divide N by 2 and round up. {internallinenumbers*} Wait, but wait. Let me check another example. Suppose N is 4. Then 4/2 is exactly 2 sheets. That makes sense. If N is 1, then 1/2 is 0.5, which rounds up to 1. So that’s correct. So the formula is simply ceiling(N / 2). But how to compute that in Python. {internallinenumbers*} In Python, if you do integer division with //, but when you need to round up, you can do (N + 1) // 2. Let’s see. For N=5: (5 +1)//2 =6//2=3. Correct. For N=4: (4+1)/2=5/2=2.5, integer division gives 2. Wait, no. Wait, 5//2 in Python is 2. So (4+1)//2 is 5//2=2. Which is correct. For N=1: (1+1)//2=1. Correct. So yes, the formula is (N +1) // 2. {internallinenumbers*} So the function is straightforward. Take the input N, compute (N +1)//2, and output it. Testing the sample input: input 5. (5+1)//2=6//2=3. Correct. Another test case: N=1 → 1. N=2 → 1. N=3 →2. All correct. So the code would be something like: Read N from input, compute (N+1)//2, print the result. So the Python code would be: n = int(input()) print( (n +1) // 2 ) {internallinenumbers*} But wait, perhaps there’s an edge case. Let’s see. For N=0? But according to constraints, N is at least 1, so no problem. {internallinenumbers*} Yes, so that’s the solution. The code is simple. Just read N, calculate (N +1)//2, print. <|end_of_thought|> Distillation solution <|begin_of_solution|> {internallinenumbers*} To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach {internallinenumbers*} The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation {internallinenumbers*} 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. {internallinenumbers*} 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. {internallinenumbers*} 3. **Output**: The result is printed, which is the minimum number of sheets needed. {internallinenumbers*} This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>
@@
Figure 10: Full formatted distillation data sample using the raw reasoning trace from our 7B RLT.
@@
Thought tokens KL reward ablation distillation example Masked distillation input <|im_start|>system <|im_start|>system Your role as an assistant involves thoroughly exploring questions... <|im_start|>user Generate an executable Python function generated from the given prompt... <|im_end|> <|im_start|>assistant Distillation reasoning trace <|begin_of_thought|> {internallinenumbers*} To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach {internallinenumbers*} The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation {internallinenumbers*} 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. {internallinenumbers*} 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. {internallinenumbers*} 3. **Output**: The result is printed, which is the minimum number of sheets needed. {internallinenumbers*} This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_thought|> Distillation solution <|begin_of_solution|> {internallinenumbers*} To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach {internallinenumbers*} The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation {internallinenumbers*} 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. {internallinenumbers*} 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. {internallinenumbers*} 3. **Output**: The result is printed, which is the minimum number of sheets needed. {internallinenumbers*} This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>
@@
Min/max reduction reward ablation distillation example Distillation reasoning trace <|begin_of_thought|> {internallinenumbers*} Okay, I need to solve this problem where Takahashi wants to print N pages double-sided, with each sheet holding two pages. The task is to find the minimum number of sheets required. Let’s think through this step by step. {internallinenumbers*} So, the problem is about packing pages into sheets. Each sheet can have two pages. But since it’s double-sided, maybe the way the pages are arranged affects how they fit. Wait, no, double-sided printing just means that two pages are printed on the same sheet, front and back. So for example, if I have page 1 on the front and page 2 on the back, that’s one sheet. But if there’s an odd number of pages, the last sheet will have only one page. For example, if N is 5, then the sample input shows 3 sheets. Let’s see: 1-2 on first sheet, 3-4 on second, 5 on third. So each sheet except possibly the last can have two pages. {internallinenumbers*} Hmm, so the approach is to calculate how many pairs we can make. For N pages, the number of pairs is N//2. But if N is odd, there’s an extra page. So the total number of sheets would be (N // 2) + (1 if N % Wait, but wait. Let’s check another example. If N is 4. Then 4//2 is 2, and 4% Another example: N=3. 3//2 is 1, 3% What about N=1? Then 1//2 is 0, 1% {internallinenumbers*} Wait, the problem says each sheet can hold two pages. So if there’s an odd number of pages, the last sheet has one page. But the problem’s sample input is 5, which works out to 3 sheets. So maybe when N is odd, the formula is (N+1)//2. Wait, 5+1 is 6, 6//2 is 3. Which is correct. For N=3, 3+1=4, 4//2=2. Correct. For N=1, 1+1=2, 2//2=1. So that’s correct. Wait, 1+1=2, 2//2=1. So the formula is (N + 1) // 2. Let’s test that. {internallinenumbers*} Wait, for N=1: (1+1)//2 = 1. Which is correct. For N=2: (2+1)/2=1.5, integer division is 1. Which is correct. For N=3: (3+1)/2=2. Correct. N=4: 5//2=2.5→2. Correct. N=5: 3. Correct. So this formula seems to work. {internallinenumbers*} Wait, but wait. Let’s see. For N=1, if we use (N+1)//2, that’s (1+1)//2=1. So yes. So the formula is (N +1)//2. But wait, in the problem statement, the sample input 5 gives output 3, which is (5+1)//2=3. So that’s correct. So the formula is (N +1)//2. Because adding 1 gives even number of pages, which can be divided by 2, and then integer division by 2 gives the correct number of sheets. {internallinenumbers*} Wait, but let me think again. For N=1, the formula gives 1. So that’s correct. For N=2, (2+1)//2=1.5→1. Which is correct. So yes, this formula works. {internallinenumbers*} Wait, but why would (N+1)//2 be the correct formula? Because when you have N pages, each sheet holds two pages, except possibly the last one. So the total number of sheets is the number of pairs (each pair is two pages) plus one if there’s an odd number. So the number of pairs is N//2, and if N is odd, there’s an extra sheet. So the total is (N//2) + (1 if N % {internallinenumbers*} Yes. Because for N even, (N+1)/2 is (N/2) +0.5 → but integer division would be N/2. For example, N=4: (4+1)/2=2.5→2. Which is correct. For N=5: (5+1)/2=3. So that’s correct. So the formula is (N +1) // 2. {internallinenumbers*} Wait, but wait. Let’s test N=1. (1+1)/2=1. Correct. N=2: (2+1)/2=1.5→1. Correct. N=3: (3+1)/2=2. Correct. N=4: 2. Correct. N=5:3. Correct. So this formula works. So the problem reduces to calculating (N + 1) // 2. {internallinenumbers*} But wait, the input is N, which is between 1 and 100. So the function needs to read N from standard input, compute (N+1)//2, and output that. {internallinenumbers*} Wait, but wait. Let’s take N=1. The sample input is 5, which gives output 3. For N=1, the output would be 1. But according to the problem statement, the minimal number of sheets. So for N=1, you need 1 sheet. Because even though you can’t fill a full sheet, you just have to use one sheet. So the formula works. So the code is straightforward. Read N, compute (N+1)//2, print that. {internallinenumbers*} Wait, but let me check another example. Let’s say N=6. (6+1)//2=3.5→3. Wait, 6//2=3. Correct. So that’s correct. {internallinenumbers*} Wait, but wait. Wait, no. Wait, (6+1)//2 is 7//2=3. Wait, 7 divided by 2 is 3.5, integer division is 3. So (6+1)//2 is 3.5 →3. So yes. So 6 pages would require 3 sheets. Each sheet has two pages. So 1-2, 3-4, 5-6. So three sheets. Correct. {internallinenumbers*} So the code is as simple as reading N, then compute the ceiling of N divided by 2. Which is (N +1)//2. So the Python function would be: Read N from input. Then compute (N +1) // 2. Print that result. Wait, but wait. Let’s test N=1 again. (1+1)//2=1. Correct. Yes. So the code is straightforward. So the code would be: Read N as integer from input. Then print (N + 1) // 2. {internallinenumbers*} But wait, wait. Let’s check N=1. (1+1)//2=1. Correct. For N=2: (2+1)//2=1. So 1 sheet. Which is correct. For N=3: 2 sheets. Correct. For N=4: 2. Correct. For N=5:3. Correct. For N=6:3. Correct. So the code is correct. Therefore, the Python function is: Read N from input, then compute (N +1) // 2 and print it. So the code would be: n = int(input()) print((n + 1) // 2) {internallinenumbers*} Wait, but wait. Let me think again. For N=1, the answer is 1. For N=2, 1. So that’s correct. So the code is correct. {internallinenumbers*} But wait, the problem says that N is the number of pages. So that’s correct. So the code is as simple as that. I think that’s it. So the answer is to compute (N +1) // 2. <|end_of_thought|> Distillation solution <|begin_of_solution|> {internallinenumbers*} To solve this problem, we need to determine the minimum number of sheets of paper required to print a document with N pages using double-sided printing, where each sheet can hold two pages. ### Approach {internallinenumbers*} The key insight here is that each sheet of paper can hold two pages. Therefore, the minimum number of sheets required is determined by dividing the number of pages by 2 and rounding up to the nearest integer. This can be efficiently computed using integer arithmetic by adding 1 to the number of pages before performing integer division by 2. This approach works because it effectively rounds up the result of dividing N by 2. ### Solution Code ```python n = int(input()) print((n + 1) // 2) ``` ### Explanation {internallinenumbers*} 1. **Reading Input**: The input is read as an integer N, which represents the number of pages. {internallinenumbers*} 2. **Calculation**: The formula `(N + 1) // 2` is used to compute the minimum number of sheets. Adding 1 to N before division ensures that any remainder from the division results in rounding up to the next whole number. {internallinenumbers*} 3. **Output**: The result is printed, which is the minimum number of sheets needed. {internallinenumbers*} This approach efficiently computes the result in constant time, O(1), making it very efficient even for the upper constraint limit. <|end_of_solution|><|im_end|>